p(atm) Pcfo - 300 V(cm) 100

If the gas's initial temperature is 730 ∘C, what is its final temperature?

If the constant pressure for this gas is pc = 5.0 atm , how many moles of gas are there?
Answer #1

Pram) Given that at initial point T=fF30+273 )K T = 1003 * P=satm P = 5x1.01 XI Nim2 p= 5.05X0SN/m2 V = 300 cm V: 300x106m2 H